OpenStudy (anonymous):

if i= (sqrt)-1, then which expression is equal to 9i(13i) A. -117 B. 117i C. 117 D. -117i

Nnesha (nnesha):

distribute parentheses by outside term

OpenStudy (anonymous):

so do you equal them like: 9i(13i)=(sqrt)-1

Nnesha (nnesha):

\[\huge\rm i = \sqrt{-1} \] just i equal sqrt{-1} multiply 9i(13i) it's not going to = sqrt-1}

OpenStudy (anonymous):

im confused...

Nnesha (nnesha):

okay first multiply 9i(13i) = ?

OpenStudy (anonymous):

117i

Nnesha (nnesha):

or first you can replace i by sqrt -1

Nnesha (nnesha):

i times i = ??

OpenStudy (anonymous):

i^2

Nnesha (nnesha):

9 i times 13 i 9 times 13 and i times i

Nnesha (nnesha):

yes right i^2 = -1

OpenStudy (anonymous):

117i^2 is the answer??

Nnesha (nnesha):

\[\huge\rm 9\sqrt{-1}(13\sqrt{-1})\] okay if we replace i by sqrt{-1} right now multiply what would u get when you multiply \[\huge\rm \sqrt{-1} \times \sqrt{-1}= ?\]

Nnesha (nnesha):

\[\huge\rm i^2 = i \times i \] \[\large\color{red}{\rm i = \sqrt{-1}}\] because i equal sqrt of -1 you can substitute sqrt -1 for i when you multiply sqrt{-1} by sqrt{-1} both square would cancel each other out so \[\sqrt{-1} \times \sqrt{-1} = -1\] so that's why i^2 = -1

OpenStudy (anonymous):

okay.

Nnesha (nnesha):

so replace i^2 bu -1

Nnesha (nnesha):

y*

OpenStudy (anonymous):

so that answer is 117i

Nnesha (nnesha):

i^2 = ???

OpenStudy (anonymous):

-1

Nnesha (nnesha):

yes so substitute -1

OpenStudy (anonymous):

-117

Nnesha (nnesha):

yes

OpenStudy (anonymous):

okay thanks alot!!!!

Nnesha (nnesha):

my pleasure :-)